# Analysis of static and dynamic batching algorithms for graph neural networks

### v1

@@ Abstract
Graph neural networks (GNN) have shown promising results for several domains such as materials science, chemistry, and the social sciences. GNN models often contain millions of parameters, and like other neural network (NN) models, are often fed only a fraction of the graphs that make up the training dataset in batches to update model parameters. The effect of batching algorithms on training time and model performance has been thoroughly explored for NNs but not yet for GNNs. We analyze two different batching algorithms for graph based models, namely static and dynamic batching. We use the Jraph library built on JAX to perform our experiments, where we compare the two batching methods for two datasets, the QM9 dataset of small molecules and the AFLOW materials database. Our experiments show that significant training time savings can be found from changing the batching algorithm, but the fastest algorithm depends on the data, model, batch size and number of training steps run. Experiments show no significant difference in model learning between the algorithms.
@@ 1 Introduction
  • The performance of the algorithm is dependent on the graph distribution in the dataset, model used, batch size and the number of training steps run.
  • Our results show that when running enough training steps to converge the model, either the static algorithm with padding to the nearest power of two or padding to the nearest multiple of 64 is fastest in terms of mean training time.
  • Across the experiments, we observe at most a 33% speedup in mean time per training step when switching from the slowest algorithm to the fastest.
  • The experiments show no significant difference in model learning between the algorithms.
@@ 4 Datasets
The datasets are not provided in a graph format. To convert them to graphs, each atom is represented as a node in the graph. For QM9 the edges are added by fully connecting the nodes. For AFLOW, the edges are added with the K-nearest neighbor algorithm based on pairwise distance and setting K equal to 24. This KNN value was chosen from CV studies in the literature (Jørgensen et al., 2018). From the AFLOW dataset, we remove duplicates and keep only calculations that contain both the enthalpy and the band structure, following the procedure in (Bechtel et al., 2023). We visualize the variety in the graph structures in the datasets in Fig. 1. From the figure we can see that the QM9 dataset node distribution appears Gaussian and centered around 17 nodes per graph. AFLOW’s node distribution has long tails despite the mean of the distribution being smaller than QM9. The distribution of edges for AFLOW and QM9 are reflective of the distrubiton of the nodes, this is because the number of edges are dependent on the number of nodes when using the KNN algorithm or fully connecting the graphs.
@@ 7 Models
We evaluate two models to get insight into any model dependent effects on our profiling results. The models take in a graph, $G$, composed of nodes (or vertices) and edges $G(n,e)$. The nodes are represented by feature vectors $h_{n}^{t}$, where the subscript represents the specific node $n$, and the superscript $t$ represents the number of times the vector has been updated.
@@
These two models were chosen, since the SchNet model is often used as a benchmark, and the MPEU has shown better results on the AFLOW dataset in a benchmark study (Bechtel et al., 2023). These models vary in size. The SchNet model has 84,768 trainable parameters, while the MPEU model contains roughly 30 times more, namely 2,553,472. Large transformer based models were avoided due to computing time constraints on GPU.
@@ 8 Profiling results
To understand how the batching algorithms work in practice, we analyze the statistics of the number of nodes in a batch before padding occurs. The histograms for both batching algorithms are shown for the QM9 dataset in Fig. 2 for 10,000 batches with a batch size of 32. As stated above, the dynamic batching algorithm checks the node and edge padding target, as well as the number of graphs already in the batch, before adding a graph to the batch. For this dataset and batch size, the dynamic algorithm’s node padding target is 576. We can see that all dynamic batches have fewer than 576 nodes before padding. This results in a one-sided distribution, roughly a truncated Gaussian distribution, where the right-hand side is cutoff near the budget. For static batching, however, no such check exists, and we observe a roughly Gaussian distribution of nodes. Similar effects are seen in the figure for the number of edges before padding.
@@
We perform timing experiments on both algorithms for four batch sizes (16, 32, 64, 128) using both datasets and models for two million training steps. This number of training steps was chosen since it resulted in converged models  (Bechtel et al., 2023). We run each combination of batch size, algorithm, model, and dataset ten times to limit noise effects from the hardware in the profiling results. Each experiment reports the mean batching time, mean gradient update step time, and the sum of the two (combined time).
@@
According to the figure, the static-$2^{N}$ and static-64 algorithms perform batching at near identical speeds. The dynamic batching, however, is clearly slower. All three algorithms show roughly linear scaling of the batching time with respect to the batch size, however, dynamic batching scales worse for larger batch sizes (i.e., has a steeper slope). This reflects the expectation that the time complexity of the batching algorithms is governed by the batch size. The dynamic algorithm is slower due to the added overhead from the bookkeeping of the padding targets. The batching results depend slightly on the dataset, but the rankings do not change and the scaling appears similar. The results for the MPEU model are shown in Fig. 8 in the Appendix.
@@
The timing results of the gradient-update step at small batch sizes is similar for all algorithms. For batch size, 128, the dynamic batching is fastest for both datasets. The static-64 and dynamic algorithm show roughly linear scaling in the update time with the batch size. The static-$2^{N}$ algorithm shows poor relative performance especially for larger batch sizes. More work needs to be done to understand the static-$2^{N}$ scaling behavior. Similar results are seen for the MPEU model in the Appendix, with the gradient-update step being slower for all algorithms for the larger parameter count MPEU model.
@@
The sum of the batching and gradient-update step mean timings, i.e., the mean combined time, roughly gives us the mean time per training step. For the SchNet model, the static-$2^{N}$ model performs the combined steps the fastest for the AFLOW dataset and for the batch size 16 in the QM9 dataset. For the batch size 64 and 128, the static-64 algorithm runs fastest. For the MPEU model the static-$2^{N}$ algorithm is fastest for batch sizes 16, 32 and 64 on AFLOW but the static-$64$ algorithm is faster for the batch size 128 on AFLOW and for all batch sizes on QM9 data. The poor update step scaling of the static-$2^{N}$ algorithm becomes the limiting factor for larger batch sizes on the QM9 dataset. The difference between the algorithms is at most 2.9 ms for the QM9 dataset using the MPEU model, from the static-$2^{N}$ to the static-64 algorithm. This difference represents a 33% speedup in training time.
@@
For fewer training steps, the number of recompilations required in the gradient-update step is the dominant factor defining the ranking of the algorithms. This number of recompilations is shown in Fig. 4 for the MPEU model on QM9 data for two million training steps. We see that the static-64 algorithm performs several orders of magnitude more recompilations than the static-$2^{N}$ and dynamic algorithms. This is to be expected, since the dynamic algorithm compiles only once, and if it encounters a graph, which is bigger than the budget, the program terminates. Moreover, for static-64 batching, JAX recompiles the gradient-update step function every time a new nearest multiple of 64 is encountered when padding, and for static-$2^{N}$ batching every time a new nearest power of two is encountered, which is less likely for large powers of two. Note that most of the recompilations for the QM9 dataset happen in the first hundred thousand training steps. For example, for the batch size 32, the static-$2^{N}$ algorithm recompiles four times while the static-64 algorithm recompiles 89 times. From training step one hundred thousand to two million, the static-$2^{N}$ algorithm does not recompile again and the static-64 algorithm recompiles only an additional twenty times. The effect of recompilations is seen in Fig. 7 and Fig. 6 in the Appendix which show the mean and median combined times after running only one hundred thousand training steps. In the figures, the static-64 algorithm has the lowest median update times while the static-$2^{N}$ has the lowest mean update times. The mean is affected by outliers while the median is not. This tells us that when removing the effect of outliers, i.e., slow gradient-update steps which are caused by recompilations, the static-64 algorithm is fastest to update the gradients. Further, this tells us that the algorithm training time rankings are dependent on the dataset (whose variability in graph sizes cause recompilations) and the of number training steps run.
@@
Figure 4: Number of recompilations on the QM9 dataset after two million training steps in the gradient-update step as a function of batch size for different batching algorithms.
@@
The effect of the algorithm on learning can be discerned by training each model to convergence. The test loss (RMSE) on the test data for both models on the QM9 data is shown in the Appendix Fig. 12. The test loss curves as a function of training time are very similar for the dynamic and static batching algorithms. The same is true for the AFLOW data which can be seen in the accompanying code repository. In our experiments the batching methods do not significantly affect the test metrics despite the fact that the dynamic algorithm often has fewer true training graphs in the batch.
Finally, the static-constant algorithm was also evaluated for a subset of data, models and batch sizes. The results are shown in the Appendix Section F. The static-constant model performs poorly compared to the other static algorithms, especially for the AFLOW dataset. This is because the AFLOW dataset contains a long tailed distribution in terms of the node/edges and as a result the static-constant model uses a large padding target that slows down the timing results. As a result, we did not employ this method for the full set of data combinations in our timing experiments.
@@ 9 Discussion
Our results show that the batching is slowest for the dynamic algorithm. This is due to the added overhead from the bookkeeping of the padding targets. All of the algorithms show linear scaling in the batching time with respect to the batch size. The static-64 and dynamic algorithms show linear scaling in the gradient-update step, while the static-$2^{N}$ algorithm shows exponential scaling. This is likely due to the exponentially larger memory sizes required when using power of two for padding. In our experiments, the static-64 or the static-$2N$ algorithms are the fastest in terms of mean training time. Which of the two is fastest, depends on the batch size, model, number of training steps run, and dataset. Our experiments show that for some datasets and batch sizes, the speedup in training time from the slowest algorithm to the fastest can be over 33%. The findings in this paper recommend that for each new dataset, model, and batch size, the user should explore possible savings offered by switching the batching algorithm.
The learning curves from the static and dynamic algorithms do not differ significantly for the two datasets examined here. This is despite the fact that the dynamic batching algorithm contains, on average, fewer labeled graphs per batch. Lastly, although the dynamic batching method does not yield the fastest training time results on GPU for our results, it’s possible for other datasets, where the graphs in the data vary significantly in size, it outperforms the static methods.
We hope that this work serves to aide users to reduce the computational cost of training graph based models. Further work should target a larger number of datasets (node-level targets as well), sets of models, and compute devices and analyze the effect of using a PyTorch based implementation. Lastly, this work could be extended to combine the batching algorithms with batching schemes that partition large graph networks across GPUs.
@@
The software to perform the batching, profiling experiments, parsing of experiments, and plotting is found in https://github.com/speckhard/icml_gnn_batching.

### v2

@@ Abstract
Graph neural networks (GNN) have shown promising results for several domains such as materials science, chemistry, and the social sciences. GNN models often contain millions of parameters, and like other neural network (NN) models, are often fed only a fraction of the graphs that make up the training dataset in batches to update model parameters. The effect of batching algorithms on training time and model performance has been thoroughly explored for NNs but not yet for GNNs. We analyze two different batching algorithms for graph based models, namely static and dynamic batching for two datasets, the QM9 dataset of small molecules and the AFLOW materials database. Our experiments show that changing the batching algorithm can provide up to a 2.7x speedup, but the fastest algorithm depends on the data, model, batch size, hardware, and number of training steps run. Experiments show that for a select number of combinations of batch size, dataset, and model, significant differences in model learning metrics are observed between static and dynamic batching algorithms.
@@ 1 Introduction
  • The performance of the algorithm is dependent on the graph distribution in the dataset, model used, batch size, hardware, and the number of training steps run.
  • Across the experiments, we observe at most a 2.7x speedup in mean time per training step when switching from the slowest algorithm to the fastest.
  • The majority of experiments show no significant difference in model learning between the algorithms, but for some combinations of the dataset, model and batch size, the dynamic batching algorithm yields significantly different test metrics than the static algorithms.
@@ 4 Datasets
The datasets are not provided in a graph format. To convert them to graphs, each atom is represented as a node in the graph. For QM9 the edges are added by fully connecting the nodes. For AFLOW, the edges are added with the K-nearest neighbor algorithm based on pairwise distance and setting K equal to 24. This KNN value was chosen from CV studies in the literature (Jørgensen et al., 2018). From the AFLOW dataset, we remove duplicates and keep only calculations that contain both the enthalpy and the band structure, following the procedure in (Bechtel et al., 2023). We visualize the variety in the graph structures in the datasets in Fig. 1. From the figure we can see that the QM9 dataset node distribution appears Gaussian and centered around 17 nodes per graph. AFLOW’s node distribution has long tails despite the mean of the distribution being smaller than QM9. The distribution of edges for AFLOW and QM9 are reflective of the distribution of the nodes, this is because the number of edges are dependent on the number of nodes when using the KNN algorithm or fully connecting the graphs.
@@ 7 Models
We evaluate three models to get insight into any model dependent effects on our profiling results. The models take in a graph, $G$, composed of nodes (or vertices) and edges $G(n,e)$. The nodes are represented by feature vectors $h_{n}^{t}$, where the subscript represents the specific node $n$, and the superscript $t$ represents the number of times the vector has been updated.
@@
Finally, we evaluate the algorithms on the equivariant PaiNN model (Schütt et al., 2021). This model differs from the MPEU and SchNet by learning not only scalar features, $s_{i}\in\mathbb{R}^{Fx1}$, for the representation of each atom but also equivariant features, $\vec{v_{i}}\in\mathbb{R}^{Fx3}$. The two representations are used to update each other. The scalar representation is updated with the following equation:
$\Delta s_{i}^{u}=a_{ss}(s_{i},\lVert V\vec{v}\rVert)+a_{sv}(s_{i},\lVert V\vec{v}\rVert)\langle U\vec{v},V\vec{v}\rangle,$ (8)
where $a_{ss}$ and $a_{sv}$ represent neural networks with non-linear activation functions and $V$ and $U$ weight matrices that need to be learned. The update function for the equivariant representation is:
$\Delta\vec{v_{i}^{u}}=a_{vv}(s_{i},\lVert V\vec{v}\rVert)U\vec{v},$ (9)
The three models were chosen, since the SchNet model is often used as a benchmark, the MPEU has shown better results on the AFLOW dataset in a benchmark study (Bechtel et al., 2023), and the PaiNN model is widely used (Kubečka et al., 2024). These models vary in size. The SchNet model is the smallest with 84,768 trainable parameters. The next largest is the PaiNN model with 177,568 trainable parameters, and the MPEU model contains 2,553,472.
@@ 8 Profiling results
To understand how the batching algorithms work in practice, we analyze the statistics of the number of nodes in a batch before padding occurs. The histograms for both batching algorithms are shown for the QM9 dataset in Fig. 2 for 10,000 batches with a batch size of 32. As stated above, the dynamic batching algorithm checks the node and edge padding target, as well as the number of graphs already in the batch, before adding a graph to the batch. For this dataset and batch size, the dynamic algorithm’s node padding target is 576. We can see that all dynamic batches have fewer than 576 nodes before padding. This results in a one-sided distribution, roughly a truncated Gaussian distribution, where the right-hand side is cutoff near the budget. In the dynamic case, the distribution of edges before padding shows an even clearer cutoff. This is due to the fact that, more often, the budget for edges is reached sooner than the budget for nodes. For static batching, however, no such check exists, and we observe a roughly Gaussian distribution of nodes. Similar effects are seen in the figure for the number of edges before padding.
@@
We perform timing experiments on both algorithms for four batch sizes (16, 32, 64, 128) using both datasets and the three models for two million training steps. This number of training steps was chosen since it resulted in converged models  (Bechtel et al., 2023). We run each combination of batch size, algorithm, model, and dataset ten times to limit noise effects from the hardware in the profiling results. Each experiment reports the mean batching time, mean gradient update step time, and the sum of the two (combined time). Below, we discuss the results on GPU, while results on CPU are found in the Appendix.
@@
According to Fig. 3, for the SchNet model, the static-$2^{N}$ and static-64 algorithms perform batching at near identical speeds. The dynamic batching, however, is clearly slower. All three algorithms show roughly linear scaling of the batching time with respect to the batch size, however, dynamic batching scales worse for larger batch sizes (i.e., has a steeper slope). This reflects the expectation that the time complexity of the batching algorithms is governed by the batch size. The dynamic algorithm is slower due to the added overhead from the bookkeeping of the padding targets. The batching results depend slightly on the dataset, but the rankings do not change and the scaling appears similar. The results for the PaiNN model are shown in Fig. 4, while the results for the MPEU model are shown in Fig. 10 in the Appendix.
Figure 4: Mean batching time (upper row), mean gradient-update time (middle row), and mean combined time (bottom row) for varying batch sizes on the AFLOW (left) and QM9 (right) data using the PaiNN model. For each datapoint, ten iterations of two million training steps are run.
@@
For the SchNet and MPEU models, the timing results of the gradient-update step at small batch sizes are similar for all algorithms. For the PaiNN model, the static-$2^{N}$ algorithm has a slower gradient update for all batch sizes. For larger batch sizes, across all models and datasets, the dynamic batching algorithm performs the gradient update the fastest. The dynamic batching algorithm is likely the fastest, since the gradient update code is never recompiled. As seen in the histograms in Fig. 1, the number of nodes in a batch before padding is typically smaller for the dynamic algorithm. This is also true after padding. GNNs like the models we examine here, typically have a node update equation, which means that inference, and therefore the gradient update step, scales roughly linearly with the number of nodes. When the difference between the number of nodes is small between the dynamic and static-64 algorithm we do not expect much difference in the gradient update step time. The static-$2^{N}$ algorithm, however, typically adds a large number of padded nodes, which explains why the algorithm is typically slower than the dynamic algorithm in the gradient update step. The static-64 and dynamic algorithm show roughly linear scaling in the update time with the batch size. The static-$2^{N}$ algorithm shows poor relative performance especially for larger batch sizes. More work needs to be done to understand the static-$2^{N}$ scaling behavior.
@@
The sum of the batching and gradient-update step mean timings, i.e., the mean combined time, roughly gives us the mean time per training step. For the SchNet model, the static-$2^{N}$ model performs the combined steps the fastest for the AFLOW dataset and for the batch size 16 in the QM9 dataset. For the batch size 64 and 128, the static-64 algorithm runs fastest. For the MPEU model the static-$2^{N}$ algorithm is fastest for batch sizes 16, 32 and 64 on AFLOW but the static-$64$ algorithm is faster for the batch size 128 on AFLOW and for all batch sizes on QM9 data. The poor update step scaling of the static-$2^{N}$ algorithm becomes the limiting factor for larger batch sizes on the QM9 dataset. For the PaiNN model, the dynamic algorithm has the shortest combined time for the AFLOW dataset and has roughly the same time requirements as the static-64 algorithm for the QM9 dataset despite having the slowest batching performance. Since the PaiNN model architecture requires a much larger update time the relative savings of the dynamic algorithm in the gradient step outweigh the poorer batching performance.
We visualize the speedup from switching algorithms for the PaiNN model on the QM9 and AFLOW datasets in Fig. 5. The speedup is defined as the ratio between the slowest algorithm combined time per step, $T_{s}$, and the faster algorithm combined time per step, $T_{x}$, as $\frac{T_{s}}{T_{x}}$ (Hennessy & Patterson, 2011). For the PaiNN model and batch size 128, the dynamic batching and static-64 algorithm have a speedup of over 2.7 when compared to the the static-$2^{N}$ algorithm. Similarly, for the PaiNN model for AFLOW and batch sizes of 32, 64 and 128, the dynamic algorithm has a speedup of roughly 1.5x when compared to the static-$2^{N}$ algorithm. For the MPEU model on the QM9 dataset, we observe 1.5x speedup when using the static-64 algorithm compared to the static-$2N$ algorithm. These observations over different datasets and models point to the large effect that changing the batching algorithm can have on the training time.
Figure 5: Speedup when switching from the slowest algorithm in terms of combined training time per step for the PaiNN model for the batch size of 64 for the AFLOW (top) and QM9 (bottom) datasets. In this case the slowest algorithm is the static-$2^{N}$ algorithm and therefore the speedup is always one.
@@
For fewer training steps, the number of recompilations required in the gradient-update step is the dominant factor defining the ranking of the algorithms. This number of recompilations is shown in Fig. 6 for the QM9 data for two million training steps. The number of recompilations is model and hardware independent. We see that the static-64 algorithm performs several orders of magnitude more recompilations than the static-$2^{N}$ and dynamic algorithms. This is to be expected, since the dynamic algorithm compiles only once, and if it encounters a graph, which is bigger than the budget, the program terminates. Moreover, for static-64 batching, JAX recompiles the gradient-update step function every time a new nearest multiple of 64 is encountered when padding, and for static-$2^{N}$ batching every time a new nearest power of two is encountered, which is less likely for large powers of two. Note that most of the recompilations for the QM9 dataset happen in the first hundred thousand training steps. For example, for the batch size 32, the static-$2^{N}$ algorithm recompiles four times while the static-64 algorithm recompiles 89 times. From training step one hundred thousand to two million, the static-$2^{N}$ algorithm does not recompile again and the static-64 algorithm recompiles only an additional twenty times. The effect of recompilations is seen in Fig. 9 and Fig. 8 in the Appendix which show the mean and median combined times after running only one hundred thousand training steps. In the figures, the static-64 algorithm has the lowest median update times while the static-$2^{N}$ has the lowest mean update times. The mean is affected by outliers while the median is not. This tells us that when removing the effect of outliers, i.e., slow gradient-update steps which are caused by recompilations, the static-64 algorithm is fastest to update the gradients. Further, this tells us that the algorithm training time rankings are dependent on the dataset (whose variability in graph sizes cause recompilations) and the of number training steps run.
@@
Figure 6: Number of recompilations on the QM9 dataset after two million training steps in the gradient-update step as a function of batch size for different batching algorithms. Note, that these numbers are model and hardware independent.
@@
The effect of the algorithm on learning can be discerned by training each model to convergence. The test loss (RMSE) as a function of the number of training steps on the QM9 data is shown in the Appendix Fig. 14. Visually, we see general agreement across models but for some batch sizes there are significant discrepancies. To quantify the difference in test metrics between two batching methods, we perform a pairwise student t-test across the final RMSE test loss. Fig. 16 in the Appendix shows a heatmap of pairwise t-test values and Fig. 17 shows the associated p-values across batching methods for the SchNet model for the QM9 dataset and batch size 64. The complete set of plots for the other combinations of datasets, models and batch sizes can be found in the Data Availability section. One can either use the t-test statistic or the p-value to determine whether to reject the null hypothesis (Hartmann et al., 2023), i.e., that the test metrics after two million training steps from different batching methods belong to the same underlying distribution. We adopt a p-value criteria from the literature (Ross, 2017), that argues that there is little evidence that the distribution of the test metrics differ significantly if the p-values are greater than 0.10. For p-values between 0.05 to 0.10, there is moderate evidence that the distributions are different. For lower p-values the evidence is strong. Ignoring the p-values that are greater than 0.10, we can look at several interesting results. For the QM9 data, SchNet model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.033. For the QM9 data, SchNet model, and batch size 32, the static-$2^{N}$ and static-64 algorithms have p-values with the dynamic algorithm of 0.067 and 0.053, respectively. For QM9 data, MPEU model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.071. These results show that for some datasets, models, and batch sizes, the dynamic algorithm gives significantly different test metrics than the static algorithm. For the majority of combinations, however, the results are similar, which may be due to use of the Adam optimizer whose adaptive estimation of first and second moments may reduce the effect of occasionally having smaller batch sizes.
The static-constant algorithm was also evaluated for a subset of data, models and batch sizes. The results are shown in the Appendix Section F. The static-constant model performs poorly compared to the other static algorithms, especially for the AFLOW dataset. This is because the AFLOW dataset contains a long tailed distribution in terms of the node/edges and as a result the static-constant model uses a large padding target that slows down the timing results. As a result, we did not employ this method for the full set of data combinations in our timing experiments.
@@ 9 Discussion
Our results show that the timing performance of the batching algorithms is dataset, model, and batch size dependent. This is true for experiments on CPU and GPU. For some cases, choosing one algorithm versus another can result in more than a 2.7x speedup in training time. The findings show that users should, for every new dataset and model, perform an initial performance test of batching algorithms. For larger hyperparameter searches or neural architecture searches, this speedup can save a very large amount of computational resources.
Diving deeper into the timing performance, the static algorithms perform the batching step faster. This is due to the added overhead from the bookkeeping of the padding targets in the dynamic algorithm. All of the algorithms show linear scaling in the batching time with respect to the batch size. The static-64 and dynamic algorithms show linear scaling in the gradient-update step, while the static-$2^{N}$ algorithm shows exponential scaling. This is likely due to the exponentially larger memory sizes required when using powers of two for padding. The gradient-update step is performed fastest by the dynamic algorithm. This is partially because the algorithm is never recompiled but also because performing inference on padded graphs, which do not influence the loss function, is typically faster. Moreover, the the static-$2^{N}$ algorithm typically has a larger number of nodes in the batch, which slows down the gradient update step.
The learning curves from the static and dynamic algorithms are mostly similar across batching algorithms, but for some batch sizes, datasets, and models they differ significantly between static and dynamic algorithms. This is because the dynamic batching algorithm contains, on average, fewer labeled graphs per batch.
We hope that this work serves to aide users to reduce the computational cost of training graph based models. Further work might focus on predicting the fastest algorithm as a function of the dataset, batch size, and model. Finally, this work could in the future be extended to combine the batching algorithms with batching schemes that partition large graph networks across GPUs.
@@
The software to perform the batching, profiling experiments, parsing of experiments, and plotting is found in https://github.com/speckhard/gnn_batching/. The student t-test heatmaps can be found at https://drive.google.com/drive/folders/1W4Fc0a7ne0z5vUDYgNc1_YuUNcjxwX3W?usp=sharing_.